OpenStudy (anonymous):

I am having a hard time trying to rationalize the following problem please help. The square root of the square root of 2- the square root of 2 divided by the square root of 2+ the square root of 2. (if anyone knows how i can type it in word and transfer it let me know but roughly this what it looks like: √((√(2-√(2)))/(√(2+√(2))))

OpenStudy (phi):

is the question rationalize A: \[\sqrt{\frac{\sqrt{2-\sqrt{2}}}{\sqrt{2+\sqrt{2}}}}\] or B: \[\frac{\sqrt{\sqrt{2-\sqrt{2}}}}{\sqrt{2+\sqrt{2}}} \] or C: neither

OpenStudy (anonymous):

The question is to rationalize the first one Choice A

OpenStudy (phi):

First step is ignore the outer square root (for the moment. Remember it is there, but concentrate on the stuff inside) multiply top and bottom by sqrt(2+sqrt(2)) what do you get?

OpenStudy (anonymous):

on the top you would get 2 and the bottom would be 6+4sqrt(2)

OpenStudy (phi):

OK, I am ignoring the outer square root (but it is still there hovering over us) \[\frac{\sqrt{2-\sqrt{2}}}{\sqrt{2+\sqrt{2}}}\cdot \frac{\sqrt{2+\sqrt{2}}}{\sqrt{2+\sqrt{2}}}\] think of the bottom as sqrt(A)*sqrt(A). It is A (in this case a complicated A)

OpenStudy (anonymous):

top would be \[(\sqrt{2-\sqrt{2}})*(\sqrt{2+\sqrt{2}}) and the bottom would be 2 sqrt(2)

OpenStudy (anonymous):

bottom\[2+\sqrt{2}\]

OpenStudy (phi):

For the top we can use sqrt(A)*sqrt(B)= sqrt(A*B) . That means we get \[\frac{\sqrt{(2-\sqrt{2})(2+\sqrt{2})}}{2+\sqrt{2}}\]

OpenStudy (anonymous):

right

OpenStudy (phi):

Now for the top, we use (a+b)(a-b)= a^2-b^2

OpenStudy (phi):

so we get (with a square root still out there) \[\frac{\sqrt{2}}{2+\sqrt{2}}\]

OpenStudy (phi):

Agree?

OpenStudy (anonymous):

yes..im following

OpenStudy (phi):

now multiply top and bottom by 2-sqrt(2) to get rid of the radical in the denominator

OpenStudy (phi):

\[\frac{\sqrt{2}}{2+\sqrt{2}}\cdot \frac{2-\sqrt{2}}{2-\sqrt{2}}= \frac{2\sqrt{2}-2}{4-2}\]

OpenStudy (anonymous):

\[2\sqrt{2}-2/2\]

OpenStudy (phi):

That simplifies to sqrt(2)-1 now we remember the outer square root, to get the final (ugly) answer: \[ \sqrt{\sqrt{2}-1} \]

OpenStudy (anonymous):

Thank you so much. thats crazy...thank you again

OpenStudy (phi):

It is what it is.